# Bumblebee gravity: spherically-symmetric solutions away from the potential minimum

### v1

@@ Abstract
In this work, we study a vector model of spontaneous spacetime-symmetry breaking coupled to gravity: the bumblebee model. The primary focus is on static spherically symmetric solutions. Complementing previous work on black hole solutions, we study the effects on the solutions when the vector field does not lie at the minimum of its potential. We first investigate the flat spacetime limit, which can be viewed as a modified electrostatics model with a nonlinear interaction term. We study the stability of classical solutions generally and in the spherically-symmetric case. We also find that certain potentials, based on hypergeometric functions, yield a Hamiltonian bounded from below. With gravity, we solve for the spherically-symmetric metric and vector field, for a variety of choices of the potential energy functions, including ones beyond the quadratic potential like the hypergeometric potentials. Special case exact solutions are obtained showing Schwarszchild-Anti de Sitter and Reissner-Nordstrom spacetimes. We employ horizon and asymptotic analytical expansions along with numerical solutions to explore the general case with the vector field away from the potential minimum. We discover interesting features of these solutions including naked singularities, repulsive gravity, and rapidly varying gravitational field near the source. Finally, we discuss observational constraints on these spacetimes using the resulting orbital behavior.
@@ I Introduction
In the late 1980s and 1990s it became of interest to re-examine the idea of extra non-scalar fields when it was found that the ground state of string field theory may give rise to nonzero vacuum expectation values of tensor fields, thereby breaking the Lorentz symmetry of the vacuum [5]. Of central importance is the idea of a potential energy function for these fields that has a minimum for nonzero values for the tensor field, similar to the spontaneous-symmetry breaking of internal symmetries in the Standard Model of particle physics [6]. The simplest case to consider is a vector field, on which we focus our attention in this work.
@@
It has become clear of late that black holes are a fruitful testing ground for new physics beyond GR and the Standard Model [7, 8, 9, 10, 11]. First and foremost, GW detection has allowed the direct study and study of propagation and polarization effects of modified models going beyond GR [12, 13, 14, 15, 16, 17]. Also black hole imaging has now become a measurable phenomena and the detailed behavior of light rays around a black hole is of interest [18].
@@
Since the 1990s, there has been substantial work on experimental and observational searches for hypothetical spacetime-symmetry breaking [19, 20], motiviated by theoretical scenarios in which it could arise [21, 22, 23, 24, 25]. While no statistical signal for a breakdown of Lorentz, CPT, or diffeomorphism symmetry has yet been found, the “parameter” space of possible effects remains large [26, 27, 28, 12, 29]. Studies of the effects of spacetime-symmetry breaking on black holes have been explored extensively in the literature in the last decades. [^1] The literature is extensive, examples include Refs. [30, 31, 32, 33, 34, 35].
@@
A general effective field theory framework for studying spacetime-symmetry breaking is now widely used [36, 23, 24]. It includes a gravity sector with arbitrary types of symmetry breaking for gravity [29], for example, there is the widely studied Ricci coupling ${\cal L}\sim s_{\mu\nu}R^{\mu\nu}$, controlled by the coefficients $s_{\mu\nu}$ [24, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. While general dynamical terms have been countenanced for coefficients like $s_{\mu\nu}$ [56], it is still challenging to study in a systematic way, and so one is often led to special toy models.
@@
Our aim in this work is to examine a class of vector models of spontaneous spacetime symmetry breaking in the context of black hole type solutions. In particular, we are interested in the effects of the potential for the vector field $B_{\mu}$, $V=V(B_{\alpha}B^{\alpha})$ on gravitational physics, especially when the field configuration is not assumed to be at its minimum.
@@
The plan of study is the following. In Sec. II, we review the background on the vector model to be studied. First, we explore the flat spacetime limit in Sec. III, with specific solutions presented or characterized for various cases of the potential $V$. In particular, we establish some generic Hamiltonian properties, such as lower boundedness in Sec. III.1. The full field equations coupled to gravity are the subject of Sec. IV and Sec. V. Sec. IV discusses some generalities including energy theorems, a simple cosmological solution, the Bekenstein “no-hair” argument, and special cases with exact solutions.
@@
Specific approaches to solving the field equations for spherical symmetry are discussed in Sec. V. In Sec. V.1, we use analytical expansions from the horizon and the asymptotically flat region combined with numerical simulations. To contrast, we consider also non-analytic expansion in the asymptotic region in Sec. V.2. Finally, we make some remarks on measuring the parameters of the model through orbital perturbations and summarize our results in VI. Our conventions include the spacetime metric signature $+$-$+$- + + +, Greek letters are spacetime indices, Latin letters are spatial indices, and covariant derivatives are $\nabla_{\mu}$. Natural units with $\hbar=c=1$ are used.
@@ II Bumblebee gravity background
A vector model of spontaneous Lorentz violation with a potential driving the symmetry breaking and a Maxwell kinetic term was first proposed in Ref. [57], dubbed the “bumblebee model”. [^2] See Ref. [58] for an explanation of the name “bumblebee”. Another vector model, dubbed “Einstein-Aether theory”, modifies the kinetic terms in the action, and specializes the potential to be a Lagrange multiplier [59]. A plethora of research on the bumblebee model with nonminimal couplings in the context of special spacetime geometries has occurred in the last decade [60, 34, 61, 62, 63, 64, 35, 65, 66, 50, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77] and prior [78, 30, 79, 80, 81]. [^3] These references are not exhaustive; a quick search on INSPIRE for “bumblebee” reveals over $170$ results. This also includes quantum issues, such as a meaningful quantization program with potentials of different forms [82, 83, 84, 85, 86].
@@
${\cal L}=\sqrt{-g}\left[\frac{1}{2\kappa}(R+\xi_{1}B_{\alpha}B^{\alpha}R+\xi_{2}B_{\alpha}B_{\beta}R^{\alpha\beta})-\frac{1}{4}B^{\mu\nu}B_{\mu\nu}-V-B_{\mu}J^{\mu}\right]+{\cal L}_{M},$ (1)
where $B_{\mu\nu}=\partial_{\mu}B_{\nu}-\partial_{\nu}B_{\mu}$ is the field strength tensor for $B_{\mu}$, and ${\cal L}_{M}$ is the matter Lagrange density which we leave unspecified. We include a possible coupling of the $B_{\mu}$ field to a conserved current $J_{\mu}$, in the case that $B_{\mu}$ is interpreted as the electromagnetic potential. Note that one can add extra “ghost” kinetic terms for $B_{\mu}$ to this action, or others, which we do not consider in this work [59]. The potential depends on the only scalar contraction available for a vector field:
@@
Spontaneous symmetry breaking of spacetime symmetry occurs when the potential has a minimum at a nonzero value of $B_{\mu}$; thereby defining the vacuum of the associated QFT. Thus, the spacetime symmetry can be broken if
@@ III.1 Field equations, Hamiltonian, and potential energy
In the case of a massive vector field, $V^{\prime}=\frac{1}{2}\mu^{2}$ this yields $\partial_{\mu}B^{\mu}=0$, when can then be used to simplify the equations. For more general nonlinear” $V^{\prime}$ cases, this trick does not work in the same way, although the condition (6) still represents a constraint on the field components [78].
We consider here an arbitrary potential $V$ and extend previous works for this case. Before solving the equation, we review the Hamiltonian evolution, as it provides useful insight for what follows. For special choices of $V$, the classical Hamiltonian dynamics have been studied [80, 60], along with similar models [87, 88]. We adopt the standard Dirac-Hamiltonian procedure for a constrained system [89, 90] and make use of results in previous works [47]. In particular, we identify the (starting) four field components: $(B_{0},B_{j})$, and the corresponding conjugate momenta $\Pi^{\mu}=\delta L/\delta{\dot{A}}_{\mu}$. We first obtain the usual primary constraint and the equation for $\vec{\Pi}$:
@@
$\vec{\Pi}$ $=\dot{\vec{B}}-\vec{\nabla}B_{0}$
where the wavey equals sign denotes a “weak” equality, or equality only when the constraints are imposed. The base hamiltonian is ${\cal H}=\frac{1}{2}\vec{\Pi}^{2}+\frac{1}{4}B_{ij}B^{ij}+V+\vec{\Pi}\cdot\vec{\nabla}B_{0}+\Pi^{0}\dot{A}_{0}$. Ensuring the constraint is preserved, yields the secondary constraint (modified Gauss’s law):
@@
After using the Hamiltonian with the primary constraint added $+u\Pi^{0}$, and ensuring that the secondary constraint is preserved, one can solve for the Lagrange multiplier $u$. The final Hamiltonian density for the constrained system is
${\cal H}=\frac{1}{2}\vec{\Pi}^{2}+\frac{1}{4}B_{ij}B^{ij}+V+\vec{\Pi}\cdot\vec{\nabla}B_{0}+\Pi^{0}\left[\frac{\vec{\nabla}\cdot(V^{\prime}\vec{B})-V^{\prime\prime}2B_{0}\vec{B}\cdot(\vec{\Pi}+\vec{\nabla}B_{0})}{-2V^{\prime\prime}B_{0}^{2}+V^{\prime}}\right].$ (9)
@@
Assuming boundary conditions where the $B_{\mu}$ and $\Pi^{\mu}$ fall off sufficiently fast at the spatial boundary, we integrate (9) over all space, and enforce the constraints (“on shell”). This yields the Hamiltonian,
@@
where $M(\alpha,\beta,\gamma)$ are Kummar confluent hypergeometric functions [93], and $z=X/\Lambda^{2}$. Here $\Lambda$ has dimensions of mass and represents the energy scale for the theory. In this work, we examine $n\geq 3$ which yields stable potentials with timelike vacuum expectation values. We further simplify the analysis by considering $n$ to be an integer, for which closed form expressions for the hypergeometric functions exist. Note that similar results with hypergeometric potentials were also found for antisymmetric tensor models [94].
To our knowledge, none of these hypergeometric potentials have been considered before in a classical analysis paper, and not with gravity couplings. One can confirm that for a class hypergeometric functions, the Hamiltonian for such potentials is bounded from below. Note that the minimum of $V$ can in fact be negative, which is interesting for cosmology and black holes as discussed later. In figure 1 we plot the scaled dimensionless version of the potential portion of the Hamiltonian density from (11),
@@
Name Form with $X$ Lower bound on H
Proca (massive vector) $V=\tfrac{1}{2}\mu^{2}X$ Yes
@@
where $f=B_{0}/\Lambda$ and $u=||\vec{B}||/\Lambda$. For a vector $B_{\mu}$ that is timelike, $u^{2}<f^{2}$, and the limit of this expression as $f^{2}\rightarrow\infty$ is $-g$, thus bounded from below. For the lightlike case, $u^{2}=f^{2}$, and it can be shown that the $-f^{4}$ term is canceled by another term, leaving a positive definite expression. For the spacelike case, $f^{2}<u^{2}$, and in this case, it can be shown that again the term $-f^{4}$ is canceled by another term. Thus the overall behavior is bounded from below (albeit not always positive definite). A study of the extrema of this function (16) with respect to $(f,u)$, shows that they exist at the points $f=\pm\sqrt{9\pm\sqrt{57}}/2,u=0$ and $f=0,u=0$. The value of ${\cal H}_{V}/(g\Lambda^{4})$ at these extrema is $0.235$ when $f=\pm\sqrt{9-\sqrt{57}}/2$, $-1.092$ when $f=\pm\sqrt{9+\sqrt{57}}/2$, and 00 when $f=0=u$. These extrema, which are all timelike values of $B_{\mu}$ turn out to be saddle points except for the solution $B_{\mu}=0$, which is a local minimum. This analysis can be carried out for the $n>3$ potentials as well. We have checked for the same properties for the $n=4$, $n=5$ hypergeometric potentials, but a thorough study along these lines of the generic hypergeometric potentials remains open. In particular, conclusions for any $n$ may be drawn using the asymptotic properties of the hypergeometric functions [92].
@@ III.2.1 Linearized solutions
$E_{r}=-\frac{1}{r^{2}}\left[c_{1}\cos(kr)+c_{2}\sin(kr)\right]+\frac{k}{r}\left(-c_{1}\sin(kr)+c_{2}\cos(kr)\right).$ (21)
Note that wave-like $1/r$ terms show up in this modified electrostatics result. Our results here are consistent with the analysis of Ref. [79]. In that paper, it was shown that $\beta_{0}$ is non-propagating and represents the solution to an initial value constraint equation rather than a dynamical equation. However, we do not include a time-dependent vector component $B_{j}$, so the solutions differ slightly. See also the cosmology solution below for more insight on this overlap.
Given the result (21), one could pursue a program of phenomenology searching for effects of the extra terms in static situations. Indeed, it was shown long ago that massive photon modifications drastically change electrostatics and have been used to tightly constrain the photon mass (e.g., $m_{\gamma}<3\times 10^{-10}\,cm^{-1}$ [97]). On dimensional grounds, we might expect similar constraints on $k$. Also, it was shown that similar modifications can arise in short-range gravity tests, coming from generic Lorentz breaking terms in the action with higher mass dimension terms in the action of an EFT framework [98]. To pursue this work for the bumblebee model interpreted as a modified electrodynamics model, we leave for future work.

### v2

@@ Abstract
In this work, we study a vector model of spontaneous spacetime-symmetry breaking coupled to gravity: the bumblebee model. The primary focus is on static spherically symmetric solutions. Complementing previous work on black hole solutions, we study the effects on the solutions when the vector field does not lie at the minimum of its potential. We first investigate the flat spacetime limit, which can be viewed as a modified electrostatic model with a nonlinear interaction term. We study the stability of classical solutions generally and in the spherically-symmetric case. We also find that certain potentials, based on hypergeometric functions, yield a Hamiltonian bounded from below. With gravity, we solve for the spherically-symmetric metric and vector field, for a variety of choices of the potential energy functions, including ones beyond the quadratic potential like the hypergeometric potentials. Special case exact solutions are obtained showing Schwarszchild-Anti de Sitter and Reissner-Nordstrom spacetimes. We employ horizon and asymptotic analytical expansions along with numerical solutions to explore the general case with the vector field away from the potential minimum. We discover interesting features of these solutions including naked singularities, repulsive gravity, and rapidly varying gravitational field near the source. Finally, we discuss observational constraints on these spacetimes using the resulting orbital behavior.
@@ I Introduction
In the late 1980s and 1990s it became of interest to reexamine the idea of extra non-scalar fields when it was found that the ground state of string field theory may give rise to nonzero vacuum expectation values of tensor fields, thereby breaking the Lorentz symmetry of the vacuum [5]. Of central importance is the idea of a potential energy function for these fields that has a minimum for nonzero values for the tensor field, similar to the spontaneous-symmetry breaking of internal symmetries in the Standard Model of particle physics [6]. The simplest case to consider is a vector field, on which we focus our attention in this work.
@@
It has become clear of late that black holes are a fruitful testing ground for new physics beyond GR and the Standard Model [7, 8, 9, 10, 11]. First and foremost, gravitational wave detection has allowed for the direct study of propagation and polarization effects of modified models going beyond GR [12, 13, 14, 15, 16, 17]. Also black hole imaging has now become a measurable phenomenon and the detailed behavior of light rays around a black hole is of interest [18].
@@
Since the 1990s, there has been substantial work on experimental and observational searches for hypothetical spacetime-symmetry breaking [19, 20], motivated by theoretical scenarios in which it could arise [21, 22, 23, 24, 25]. While no statistically-significant signal for a breakdown of Lorentz, CPT, or diffeomorphism symmetry has yet been found, the “parameter” space of possible effects remains large [26, 27, 28, 12, 29]. Studies of the effects of spacetime-symmetry breaking on black holes have been published in the last decades. [^1] The literature is extensive, examples include Refs. [30, 31, 32, 33, 34, 35].
@@
A general effective field theory (EFT) framework for studying spacetime-symmetry breaking is now widely used [36, 23, 24]. It includes a gravity sector with arbitrary types of symmetry breaking for gravity [29], for example, there is the widely studied Ricci coupling ${\cal L}\sim s_{\mu\nu}R^{\mu\nu}$, controlled by the coefficients $s_{\mu\nu}$ [24, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. While general dynamical terms have been countenanced for coefficients like $s_{\mu\nu}$ [56], it is still challenging to study in a systematic way, and so one is often led to special toy models.
@@
Our aim in this work is to examine a class of vector models of spontaneous spacetime symmetry breaking in the context of black hole-type solutions. In particular, we are interested in the effects of the potential for the vector field $B_{\mu}$, $V=V(B_{\alpha}B^{\alpha})$ on gravitational physics, especially when the field configuration is not assumed to be at its minimum.
@@
The plan of study is the following. In Sec. II, we review the background on the vector model to be studied. First, we explore the flat spacetime limit in Sec. III, with specific solutions presented or characterized for various cases of the potential $V$. In particular, we establish some generic Hamiltonian properties, such as lower boundedness in Sec. III.1. The full field equations coupled to gravity are the subject of Sec. IV and Sec. V. Section IV discusses some generalities including energy theorems, a simple cosmological solution, the Bekenstein “no-hair” argument, and special cases with exact solutions.
@@
Specific approaches to solving the field equations for spherical symmetry are discussed in Sec. V. In Sec. V.1, we use analytical expansions from the horizon and the asymptotically flat region combined with numerical simulations. To contrast, we consider also a nonanalytic expansion in the asymptotic region in Sec. V.2. Finally, we make some remarks on measuring the parameters of the model through orbital perturbations and summarize our results in Sec. VI. Our conventions include the spacetime metric signature +++-+++, Greek letters are spacetime indices, Latin letters are spatial indices, and covariant derivatives are $\nabla_{\mu}$. Natural units with $\hbar=c=1$ are used.
@@ II Bumblebee gravity background
A vector model of spontaneous Lorentz violation with a potential driving the symmetry breaking and a Maxwell kinetic term was first proposed in Ref. [57], dubbed the “bumblebee model.” [^2] See Ref. [58] for an explanation of the name “bumblebee.” Another vector model, dubbed “Einstein-Aether theory”, modifies the kinetic terms in the action, and specializes the potential to be a Lagrange multiplier [59]. A plethora of research on the bumblebee model with nonminimal couplings in the context of special spacetime geometries has occurred in the last decade [60, 34, 61, 62, 63, 64, 35, 65, 66, 50, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77] and prior [78, 30, 79, 80, 81]. [^3] These references are not exhaustive; a quick search on INSPIRE for “bumblebee” reveals over $170$ results. This also includes quantum issues, such as a meaningful quantization program with potentials of different forms [82, 83, 84, 85, 86].
@@
${\cal L}$ $=\sqrt{-g}\Big{[}\frac{1}{2\kappa}(R+\xi_{1}B_{\alpha}B^{\alpha}R+\xi_{2}B_{\alpha}B_{\beta}R^{\alpha\beta})$ (1)
$\phantom{space}-\frac{1}{4}B^{\mu\nu}B_{\mu\nu}-V-B_{\mu}J^{\mu}\Big{]}+{\cal L}_{M},$
where $B_{\mu\nu}=\partial_{\mu}B_{\nu}-\partial_{\nu}B_{\mu}$ is the field strength tensor for $B_{\mu}$, and ${\cal L}_{M}$ is the matter Lagrange density which we leave unspecified. We include a possible coupling of the $B_{\mu}$ field to a conserved current $J_{\mu}$, in the case that $B_{\mu}$ is interpreted as the electromagnetic potential. Note that one can add extra “ghost” kinetic terms for $B_{\mu}$ to this action, or others, which we do not consider in this work. The potential depends on the only scalar contraction available for a vector field:
@@
Spontaneous symmetry breaking of spacetime symmetry occurs when the potential has a minimum at a nonzero value of $B_{\mu}$; thereby defining the vacuum of the associated Quantum Field Theory. Thus, the spacetime symmetry can be broken if
@@
Note that the condition (4) implies that this work differs from other phenomenological approaches where the effects of $\sim\xi BBR$ are the focus. The latter approaches generate background “coefficients” that can be measured like $s_{\mu\nu}\sim\xi b_{\mu}b_{\nu}$, and their presence clearly indicates symmetry breaking, at least in weak-field gravity scenarios. Nonetheless, our focus on asymptotically flat solutions will generate values of $B_{\mu}$ in the asymptotic region that can be identified as vector coefficients for Lorentz violation akin to the $b_{\mu}$ coefficients in the EFT framework [87].
@@ III.1 Field equations, Hamiltonian, and potential energy
In the case of a massive vector field, $V^{\prime}=\frac{1}{2}\mu^{2}$ this yields $\partial_{\mu}B^{\mu}=0$, which can then be used to simplify the equations. For more general nonlinear” $V^{\prime}$ cases, this trick does not work in the same way, although the condition (6) still represents a constraint on the field components [78].
We consider here an arbitrary potential $V$ and extend previous works for this case. Before solving the equation, we review the Hamiltonian evolution, as it provides useful insight for what follows. For special choices of $V$, the classical Hamiltonian dynamics have been studied [80, 60], along with similar models [88, 89]. We adopt the standard Dirac-Hamiltonian procedure for a constrained system [90, 91] and make use of results in previous works [47]. In particular, we identify the (starting) four field components: $(B_{0},B_{j})$, and the corresponding conjugate momenta $\Pi^{\mu}=\delta L/\delta{\dot{A}}_{\mu}$. The Lagrangian in this limit is $L=-\int d^{3}x(\tfrac{1}{4}B^{\mu\nu}B_{\mu\nu}+V)$. We first obtain the usual primary constraint and the equation for $\vec{\Pi}$:
@@
$\vec{\Pi}$ $=\dot{\vec{B}}-\vec{\nabla}B_{0},$
where the wavy equals sign denotes a “weak” equality, or equality only when the constraints are imposed. The base, or beginning Hamiltonian density is ${\cal H}_{0}=\pi^{\mu}\dot{B}_{\mu}-{\cal L}=\frac{1}{2}\vec{\Pi}^{2}+\frac{1}{4}B_{ij}B^{ij}+V+\vec{\Pi}\cdot\vec{\nabla}B_{0}+\Pi^{0}\dot{A}_{0}$. The next step is to check the evolution of the primary constraint $\dot{\Pi}^{0}=\{\Pi^{0},H\}$. The brackets here are Poisson brackets and explicit formulas can be found in standard references, for example, see Appendix B in Ref. [47], or the appendix in Ref. [88]. Ensuring the constraint is preserved, $\dot{\Pi^{0}}\approx 0$, yields the secondary constraint (modified Gauss’s law):
@@
After using the Hamiltonian with the primary constraint added $u\Pi^{0}$, $\cal H_{A}$, and ensuring that the secondary constraint is preserved, one can solve for the Lagrange multiplier $u$ in terms of functions of the fields $\Pi^{\mu}$ and $B_{\mu}$. The final Hamiltonian density for the constrained system is ${\cal H}_{A}$ with the formula for $u$ inserted:
${\cal H}$ $=\frac{1}{2}\vec{\Pi}^{2}+\frac{1}{4}B_{ij}B^{ij}+V+\vec{\Pi}\cdot\vec{\nabla}B_{0}$ (9)
$\phantom{sp}+\Pi^{0}\left[\frac{\vec{\nabla}\cdot(V^{\prime}\vec{B})-V^{\prime\prime}2B_{0}\vec{B}\cdot(\vec{\Pi}+\vec{\nabla}B_{0})}{-2V^{\prime\prime}B_{0}^{2}+V^{\prime}}\right].$
@@
Assuming boundary conditions where the $B_{\mu}$ and $\Pi^{\mu}$ fall off sufficiently fast at the spatial boundary, we integrate equation (9) over all space, and enforce the constraints (“on shell”). This yields the Hamiltonian,
@@
where $M(\alpha,\beta,\gamma)$ are Kummer confluent hypergeometric functions [94], and $z=X/\Lambda^{2}$. Here $\Lambda$ has dimensions of mass and represents the energy scale for the theory. In this work, we examine $n\geq 3$ which yields stable potentials with timelike vacuum expectation values. We further simplify the analysis by considering $n$ to be an integer, for which closed form expressions for the hypergeometric functions exist. Note that similar results with hypergeometric potentials were also found for antisymmetric tensor models [95].
To our knowledge, none of these hypergeometric potentials have been considered before in a classical analysis paper, and not with gravity couplings. One can confirm that for a class of hypergeometric functions, the Hamiltonian for such potentials is bounded from below. Note that the minimum of $V$ can in fact be negative, which is interesting for cosmology and black holes as discussed later. In figure 1 we plot the scaled dimensionless version of the potential portion of the Hamiltonian density from (11),
@@
Lower
bound
Name Form with $X$ on H
Proca (massive $V=\tfrac{1}{2}\mu^{2}X$ Yes
  $s$ vector)
@@
[…]
@@ III.2.1 Linearized solutions
$E_{r}$ $=-\frac{1}{r^{2}}\left[c_{1}\cos(kr)+c_{2}\sin(kr)\right]$ (21)
$\phantom{sp}+\frac{k}{r}\left[-c_{1}\sin(kr)+c_{2}\cos(kr)\right].$
Note that wave-like $1/r$ terms show up in this modified electrostatic result. Our results here are consistent with the analysis of Ref. [79]. In that paper, it was shown that $\beta_{0}$ is nonpropagating and represents the solution to an initial value constraint equation rather than a dynamical equation. However, we do not include a time-dependent vector component $B_{j}$, so the solutions differ slightly.
Given the result (21), one could pursue a program of phenomenology searching for effects of the extra terms in static situations. Indeed, it was shown long ago that massive photon modifications drastically change electrostatics and have been used to tightly constrain the photon mass (e.g., $m_{\gamma}<3\times 10^{-10}\,cm^{-1}$ [98]). On dimensional grounds, we might expect similar constraints on $k$. Also, it was shown that similar modifications can arise in short-range gravity tests, coming from generic Lorentz-breaking terms, with higher mass dimension, in the action of an EFT framework [99]. To pursue this work for the bumblebee model interpreted as a modified electrodynamics model, we leave for the future.